[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Aug 22, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Aug 23 02:36:47 UTC 2017


HSU Shift Log for Tue Aug 22 2017 22:36:38 GMT-0400 (EDT)

Shift: R

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Pablo and Marc L are working on the phone issues.  
As of now only 2 phones are operating... 4415, 4417.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170823/4606fa86/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list